# Instability and breaking of internal waves in a horizontal shear layer

### v1

@@ Abstract
The behaviour of internal waves propagating in a background shear flow is studied in the case where the direction of shear is orthogonal to gravity. Ray-tracing theory is used to predict properties of the wave state at locations where instability occurs. It is found that local wave energy growth may result from two distinct mechanisms: an increase in wave steepness due to refraction by the shear, or an increase in streamwise velocity perturbations due to wave advection of the background flow. Based on the initial conditions, a dimensionless parameter $F$ is constructed to predict the relative importance of these two mechanisms in facilitating wave-breaking according to the local ratio of perturbation kinetic to potential energy. When $F$ is small but wave amplitudes are locally large, energy remains equipartitioned and subsequent instabilities are expected to develop due to a combination of shear and convection. On the other hand, as $F$ increases, kinetic energy dominates and wave advection of momentum may instead cause breaking to become increasingly driven by enhanced vertical shear. To test these predictions, fully nonlinear direct numerical simulations are conducted, spanning a range of wave-breaking dynamics. Good qualitative agreement with the theory is found despite substantial departures from the underlying assumptions. Wave breaking leads to significant turbulent dissipation, which in some cases greatly exceeds the initial wave energy. Momentum and energy transfers between the wave, background flow and turbulence are found to be sensitive to the dynamics of breaking, as are the mixing properties.
@@ 1 Introduction
In both the atmosphere and the ocean, a significant portion of energy is stored in and transported by internal gravity wave motions. The means by which these waves break and dissipate their energy via small-scale turbulence is a classical and important problem with an extremely broad scope, encompassing a variety of linear and nonlinear interactions and instabilities over a range of scales. Wave breaking is often inferred observationally by considering one-dimensional vertical spectra of shear and isopycnal strain. At vertical scales larger than O(10m), there is generally a close match between observed spectra and the empirical Garrett-Munk spectrum modelling dynamics dominated by weakly nonlinear internal waves (garrett1975space). On the other hand, energetic turbulence typically exhibits an inertial subrange over scales smaller than the Ozmidov length scale (below which the vertical extents of turbulent motions are not affected at leading order by the stratification). It is routinely observed that the smallest scales of internal waves are larger than the largest scales of isotropic turbulence (kunze2019unified; dasaro2022internal). The dynamics across the resulting intermediate range of scales are in general not well understood, though evidence suggests they may be of leading order importance in determining the properties of turbulence and mixing at dissipative scales. For example, it is becoming increasingly apparent from observations in a variety of environments that the mixing efficiency of stratified turbulence – often assumed to be constant throughout the ocean interior – exhibits up to an order of magnitude variability over spatial and temporal scales significantly larger than those associated with turbulence (for recent examples, see alford2025buoyancy; lewin2025multiscale; ijichi2018observed). The implication is therefore that the dynamics at the scales of wave breaking are of leading order importance for the subsequent properties of mixing, a feature framed by caulfield2021layering in the terms ‘history matters’.
@@
To capture the influence of dynamics at scales over which internal waves dominate energetics and to model downscale energy transfer in a mathematically tractable way, our study will consider the evolution and eventual instability of linear monochromatic internal waves under the Boussinesq approximation and in the limit of weak rotation $f\ll N$, where $f$ and $N$ are the Coriolis and buoyancy frequency, respectively. Floquet theory may be used to show that even plane periodic waves of this kind are unstable at arbitrarily small amplitudes to a parametric instability driven by wave self-interaction (mied1976occurrence; klostermeyer1982parametric; lombard1996instability). However, the growth rate of this instability is typically expected to be small, in particular compared with the rate of modulation of the local wave amplitude due to wave-mean flow interactions (see, e.g. sutherland2006internal). In either case, when the wave amplitude is sufficiently large (such that instabilities grow quickly) or the frequency is sufficiently small, a local linear stability analysis may be performed on vertical velocity and buoyancy profiles derived from or approximating the wave polarisation relations (fritts2003gravity; dunkerton1997shear). This provides potentially useful information about, for example, the roles of convection and shear in the resulting turbulent state (lelong1998inertiaa; lelong1998inertiab), which are found to vary significantly with the properties of the background wave. In this framework, there is therefore motivation to explore the range of possible unstable configurations that are dynamically accessible during the lifecycle of an internal wave.
@@
A classical paradigm for the modulation and eventual breaking of small-scale, vertically propagating wave packets is the rapid increase in vertical wavenumber due to strong refraction near critical levels in a vertical shear flow. Such critical levels, originally studied by booker1967critical and bretherton1966propagation, are locations where the horizontal phase velocity of a wave group approaches the speed of the horizontal background flow. A Wentzel-Kramers-Brillouin-Jeffreys (WKBJ) analysis may be used to demonstrate that the intrinsic frequency (the frequency measured by an observer traveling at the speed of the background flow) decreases and approaches zero at the critical level (in the absence of rotation). The wave becomes increasingly steep leading to the possibility of breaking, as has been captured experimentally by koop1981preliminary; thorpe1981experimental and in turbulence-resolving numerical simulations by winters1994three; dornbrack1998turbulent; howland2021shear. Approximate local forms for the vertical velocity and density profiles indicate that strong vertical shear and statically unstable density profiles are both possible (winters1992instability); correspondingly, the emerging modes of instability are dynamically rich and inherently three-dimensional, with important consequences for the resulting turbulent energetics (winters1994three). A notable success of this paradigm has been the eikonal theory of ray tracing, which, when applied to test waves in a measured background shear flow, may be used to estimate the fraction of energy that is cascaded to small scales and eventually dissipated (henyey1986energy). This method forms the basis of the finescale parameterisation approach for estimating energy dissipation from measurements of the larger-scale internal wave field (polzin2014finescale).
@@
The problem of internal wave modulation by a horizontal shear has received relatively less attention in the literature. In this case, WKBJ analysis indicates the existence of a new type of critical level where the intrinsic wave frequency approaches the buoyancy frequency (olbers1981propagation; basovich1984internal; badulin1985trapping), resulting in dynamics that are characteristically distinct from the vertical shear scenario described above. These critical levels are sometimes referred to as baroclinic critical layers and have been show to play a significant role in the dynamics of rotating stratified shear flows (marcus2013three) and tilted stratified vortices (boulanger2007structure). Nonlinear dynamics within baroclinic critical levels under a steady wave forcing in a stratified uniform horizontal shear flow have been described by wang2020nonlineari; wang2020nonlinearii. More relevant to this study is the work of staquet2002transport, who investigated the transient problem of an inertia-gravity wave packet approaching a critical level in a horizontal shear layer using numerical simulations. Wave breaking and significant subsequent turbulence production was observed due to a transient growth in wave energy extracted from the shear in the vicinity of the critical level. Building on this result, bakas2009gravitya; bakas2009gravityb analysed the process by which linear internal waves may extract energy from a background unbounded uniform horizontal shear flow, thus potentially leading to instability and wave breaking. They showed that, in fact, two distinct mechanisms for energy perturbation growth exist. The first corresponds to wave amplitude growth due to extraction of energy from the background shear through downgradient Reynolds stresses, consistent with the dynamics near critical levels observed by staquet2002transport. The second results from wave advection of the mean flow momentum that may produce significant vertical shears, thus providing an alternative route to wave breaking. This latter mechanism operates in the same way as the well-known lift-up mechanism for unstratified flows (ellingsen1975stability).
@@
Motivated by the discussion above, in this study we revisit the initial value problem considered by staquet2002transport consisting of a monochromatic internal wave packet in a background horizontal shear layer
@@
where $\Delta u^{*}$ and $h^{*}$ are constants characterising the velocity and length scales of the shear and stars denote dimensional variables. The linear stability of this base flow in a uniform stratification was studied by deloncle2007three, who showed that the most unstable mode is two-dimensional with zero vertical wavenumber. However, as the (horizontal) Froude number $Fr_{h}=\Delta u^{*}/N^{*}h^{*}$ decreases, the growth rates of modes with vertical wavenumber $k_{z}>0$ increase, thereby increasing the tendency for vertical structure to emerge in practice. Indeed, the fully nonlinear simulations of basak2006dynamics highlight such three-dimensional behaviour, though turbulence was not sustained for Froude numbers $Fr_{h}\leq 1$. Recently, lewin2024evidence demonstrated the existence of energetic turbulence in this strongly stratified regime, provided an additional small-amplitude wave-like perturbation was supplied as an initial condition. Waves were able to rapidly extract energy from the horizontal shear and amplify vertical shear in the flow by multiple orders of magnitude through the lift-up mechanism described by bakas2009gravitya, resulting in a significant departure from the dynamics observed by basak2006dynamics. In this study, the linear instability of the stratified horizontal shear flow is suppressed by choosing a sufficiently large shear layer width $h^{*}$ relative to the domain size in the $x$-direction (precisely, unstable perturbations are restricted to wavelengths $\lambda_{x}^{*}>2\pi h^{*}$ (deloncle2007three)).
We investigate the local structure and amplitude of wave-like perturbations in the shear flow (1) using ray-tracing theory derived under the WKBJ approximation. A principle aim is to explore whether this method, similar to that of winters1992instability for the vertical shear case, can make useful predictions regarding the local flow fields of the waves from which instabilities may develop and lead to breaking. In particular, we wish to determine whether the effects of the two mechanisms for wave energy growth in a horizontal shear described by bakas2009gravitya may be captured by the local theory. Using this theory as a guide, we then proceed to construct a parameter space (based on the scales associated with the wave and shear flow) for investigation using fully nonlinear direct numerical simulations (DNS). In cases where wave-breaking leads to turbulence, a key question to be addressed is: how do the eventual bulk momentum and energy exchanges between the mean flow and perturbations depend on the route to wave breaking? An obvious point of reference is the vertical critical level problem, where it has been typically observed that a significant fraction of the incoming wave energy is absorbed by the mean flow near the critical level, even in cases where wave-breaking results in considerable turbulent dissipation (howland2021shear; winters1994three; thorpe1981experimental).
@@ 2.1 Initial value problem
The Reynolds number $Re_{w}\equiv N^{*}/(K^{*2}\nu^{*})$ and Prandtl number $Pr\equiv\nu^{*}/\kappa^{*}$ are defined in terms of the molecular kinematic viscosity $\nu^{*}$ and buoyancy diffusivity $\kappa^{*}$. Note that, under this choice of non-dimensionalisation, the coefficient in front of the buoyancy term in the momentum equation (3a) is unity.
@@
Figure 1: Schematic showing a representative initial condition for the problem under consideration.
@@
with amplitudes given by (5). Taking $y_{0}<0$, the sign of $k_{y0}$ is chosen to be positive such that the wave packet propagates in the positive $y$-direction. We note that wave packets of this form have also been studied by both staquet2002transport and bakas2009gravitya. In dimensionless form, the background shear flow into which the wave will propagate is
@@ 2.2 Linear dynamics
where $\epsilon$ is a small parameter characterising the relative scale of phase variations. In the problem at hand, the WKBJ framework describes waves with wavelengths that are asymptotically small relative to the width of the wave envelope $a$ and the width of shear layer $h$. Precisely, a local frequency $\omega(\mathbf{x},t)$ and wavenumber $\mathbf{k}(\mathbf{x},t)$ may be defined in terms of the phase function $\varphi(\mathbf{k})$ as
@@
where $\Omega\equiv\omega-Uk_{x}$ is the intrinsic frequency measured by an observer travelling at the speed of the background flow. The frequency $\omega$ measured by a stationary observer is often referred to as the absolute (or extrinsic) frequency. The two are related by the Doppler shift term $U(y)k_{x}$. For a nontrivial balance in (12), we must have $\Omega=O(1)$. It follows that the first two terms are $O(\epsilon^{-2})$, so that the standard dispersion relation (6) is recovered when $(\partial U/\partial y)k_{x}k_{y}=O(\epsilon^{-1})$. Because $k_{x}$ is a priori assumed $O(1/\epsilon)$, the usual approach in the literature is to assume that $U$ (and hence $\partial U/\partial y$) is $O(\epsilon)$ (see, e.g. olbers1981propagation). However, it will be shown below that, due to the fact that the background flow does not vary in the $x$ direction, $k_{x}$ is constant along rays. Therefore, the assumption that $k_{x}=1/\epsilon$ and hence $U=O(\epsilon)$ may be relaxed (provided the wave amplitude varies sufficiently slowly in $x$), as bretherton1966propagation himself pointed out. Thus we may instead consider the case $k_{x}=O(1)$ with $U=O(1)$, in which case local polarisation relations are given by

### v2

# arXiv:2511.13280v2[physics.flu-dyn] 02 Mar 2026

@@ Abstract
The behaviour of internal waves propagating in a background shear flow is studied in the case where the direction of shear is orthogonal to gravity. Ray-tracing theory is used to predict properties of the wave state at locations where instability occurs. Local wave energy growth is found to result from two distinct mechanisms: an increase in wave steepness due to refraction by the shear, or an increase in streamwise velocity perturbations due to wave advection of the background flow. Based on the initial conditions, a dimensionless perturbation energy ratio $F$ is constructed to predict the relative importance of these two mechanisms in facilitating wave-breaking. When $F$ is small and waves become locally steep, perturbation kinetic and potential energy remain approximately equipartitioned and subsequent instabilities are expected to develop due to a combination of shear and convection. On the other hand, as $F$ increases, kinetic energy dominates and wave advection of momentum may instead cause breaking to become increasingly driven by enhanced vertical shear. To test these predictions, fully nonlinear direct numerical simulations are conducted, spanning a range of wave-breaking dynamics. Good qualitative agreement with the theory is found despite substantial departures from the underlying assumptions. Wave breaking leads to significant turbulent dissipation, which in some cases greatly exceeds the initial wave energy. Momentum and energy transfers between the wave, background flow and turbulence are found to be sensitive to the dynamics of breaking, as are the mixing properties.
@@ 1 Introduction
In both the atmosphere and the ocean, a significant portion of energy is stored in and transported by internal gravity wave motions. The means by which these waves break and dissipate their energy via small-scale turbulence is a classical and important problem with an extremely broad scope, encompassing a variety of linear and nonlinear interactions and instabilities over a range of scales. Wave breaking is often inferred observationally by considering one-dimensional vertical spectra of shear and isopycnal strain. At vertical scales larger than O(10m), there is generally a close match between observed spectra and the empirical Garrett-Munk spectrum modelling dynamics dominated by weakly nonlinear internal waves (Garrett and Munk, 1975). On the other hand, energetic turbulence typically exhibits an inertial subrange over scales smaller than the Ozmidov length scale (below which the vertical extents of turbulent motions are not affected at leading order by the stratification). It is routinely observed that the smallest scales of internal waves are larger than the largest scales of isotropic turbulence (Kunze, 2019; D’Asaro, 2022). The dynamics across the resulting intermediate range of scales are in general not well understood, though evidence suggests they may be of leading order importance in determining the properties of turbulence and mixing at dissipative scales. For example, it is becoming increasingly apparent from observations in a variety of environments that the mixing efficiency of stratified turbulence – often assumed to be constant throughout the ocean interior – exhibits up to an order of magnitude variability over spatial and temporal scales significantly larger than those associated with turbulence (for recent examples, see Alford et al. (2025); Lewin et al. (2025); Ijichi and Hibiya (2018)). The implication is therefore that the dynamics at the scales of wave breaking are of leading order importance for the subsequent properties of mixing, a feature framed by Caulfield (2021) in the terms ‘history matters’.
@@
To capture the influence of dynamics at scales over which internal waves dominate energetics and to model downscale energy transfer in a mathematically tractable way, our study will consider the evolution and eventual instability of linear monochromatic internal waves under the Boussinesq approximation and in the limit of weak rotation $f\ll N$, where $f$ and $N$ are the Coriolis and buoyancy frequency, respectively. Floquet theory may be used to show that even plane periodic waves of this kind are unstable at arbitrarily small amplitudes to a parametric instability driven by wave self-interaction (Mied, 1976; Klostermeyer, 1982; Lombard and Riley, 1996a). However, the growth rate of this instability is typically expected to be small, in particular compared with the rate of modulation of the local wave amplitude due to wave-mean flow interactions (see, e.g. Sutherland, 2006a). In either case, when the wave amplitude is sufficiently large (such that instabilities grow quickly) or the frequency is sufficiently small, a local linear stability analysis may be performed on vertical velocity and buoyancy profiles derived from or approximating the wave polarisation relations (Fritts and Alexander, 2003; Dunkerton, 1997). This provides potentially useful information about, for example, the roles of convection and shear in the resulting turbulent state (Lelong and Dunkerton, 1998a, b), which are found to vary significantly with the properties of the background wave. In this framework, there is therefore motivation to explore the range of possible unstable configurations that are dynamically accessible during the lifecycle of an internal wave.
@@
A classical paradigm for the modulation and eventual breaking of small-scale, vertically propagating wave packets is the rapid increase in vertical wavenumber due to strong refraction near critical levels in a vertical shear flow. Such critical levels, originally studied by Booker and Bretherton (1967) and Bretherton (1966), are locations where the horizontal phase velocity of a wave group approaches the speed of the horizontal background flow. A Wentzel-Kramers-Brillouin-Jeffreys (WKBJ) analysis may be used to demonstrate that the intrinsic frequency (the frequency measured by an observer travelling at the speed of the background flow) decreases and approaches zero at the critical level (in the absence of rotation). The wave becomes increasingly steep leading to the possibility of breaking, as has been captured experimentally by Koop (1981); Thorpe (1981) and in turbulence-resolving numerical simulations by Winters and D’Asaro (1994); Dörnbrack (1998); Howland et al. (2021). Approximate local forms for the vertical velocity and density profiles indicate that strong vertical shear and statically unstable density profiles are both possible (Winters and Riley, 1992); correspondingly, the emerging modes of instability are dynamically rich and inherently three-dimensional, with important consequences for the resulting turbulent energetics (Winters and D’Asaro, 1994). A notable success of this paradigm has been the eikonal theory of ray tracing, which, when applied to test waves in a measured background shear flow, may be used to estimate the fraction of energy that is cascaded to small scales and eventually dissipated (Henyey et al., 1986). This method forms the basis of the finescale parameterisation approach for estimating energy dissipation from measurements of the larger-scale internal wave field (Polzin et al., 2014).
@@
The problem of internal wave modulation by a horizontal shear has received relatively less attention in the literature. In this case, WKBJ analysis indicates the existence of a new type of critical level where the intrinsic wave frequency approaches the buoyancy frequency (Olbers, 1981; Basovich and Tsimring, 1984; Badulin et al., 1985), resulting in dynamics that are characteristically distinct from the vertical shear scenario described above. These critical levels are sometimes referred to as baroclinic critical layers and have been shown to play a significant role in the dynamics of rotating stratified shear flows (Marcus et al., 2013) and tilted stratified vortices (Boulanger et al., 2007). Nonlinear dynamics within baroclinic critical levels under a steady wave forcing in a stratified uniform horizontal shear flow have been described by Wang and Balmforth (2020, 2021). More relevant to this study is the work of Staquet and Huerre (2002), who investigated the transient problem of an inertia-gravity wave packet approaching a critical level in a horizontal shear layer using numerical simulations. Wave breaking and significant subsequent turbulence production was observed due to a transient growth in wave energy extracted from the shear in the vicinity of the critical level. Building on this result, Bakas and Farrell (2009a, b) analysed the process by which linear internal waves may extract energy from a background unbounded uniform horizontal shear flow, thus potentially leading to instability and wave breaking. They showed that, in fact, two distinct mechanisms for energy perturbation growth exist. The first corresponds to wave amplitude growth due to extraction of energy from the background shear through downgradient Reynolds stresses, consistent with the dynamics near critical levels observed by Staquet and Huerre (2002). The second results from wave advection of the mean flow momentum that may produce significant vertical shears, thus providing an alternative route to wave breaking. This latter mechanism operates in the same way as the well-known lift-up mechanism for unstratified flows (Ellingsen and Palm, 1975).
@@
Motivated by the discussion above, in this study we revisit the initial value problem considered by Staquet and Huerre (2002) consisting of a monochromatic internal wave packet in a background horizontal shear layer
@@
where $\Delta u^{*}$ and $h^{*}$ are constants characterising the velocity and length scales of the shear and stars denote dimensional variables. The linear stability of this base flow in a uniform stratification was studied by Deloncle et al. (2007), who showed that the most unstable mode is two-dimensional with zero vertical wavenumber. However, as the (horizontal) Froude number $Fr_{h}=\Delta u^{*}/N^{*}h^{*}$ decreases, the growth rates of modes with vertical wavenumber $k_{z}>0$ increase, thereby increasing the tendency for vertical structure to emerge in practice. Indeed, the fully nonlinear simulations of Basak and Sarkar (2006) highlight such three-dimensional behaviour, though turbulence was not sustained for Froude numbers $Fr_{h}\leq 1$. Recently, Lewin and Caulfield (2024) demonstrated the existence of energetic turbulence in this strongly stratified regime, provided an additional small-amplitude wave-like perturbation was supplied as an initial condition. Waves were able to rapidly extract energy from the horizontal shear and amplify vertical shear in the flow by multiple orders of magnitude through the lift-up mechanism described by Bakas and Farrell (2009a), resulting in a significant departure from the dynamics observed by Basak and Sarkar (2006). In this study, the linear instability of the stratified horizontal shear flow is suppressed by choosing a sufficiently large shear layer width $h^{*}$ relative to the domain size in the $x$-direction (precisely, unstable perturbations are restricted to wavelengths $\lambda_{x}^{*}>2\pi h^{*}$ (Deloncle et al., 2007)).
We investigate the local structure and amplitude of wave-like perturbations in the shear flow (1) using ray-tracing theory derived under the WKBJ approximation. A principle aim is to explore whether this method, similar to that of Winters and Riley (1992) for the vertical shear case, can make useful predictions regarding the local flow fields of the waves from which instabilities may develop and lead to breaking. In particular, we wish to determine whether the effects of the two mechanisms for wave energy growth in a horizontal shear described by Bakas and Farrell (2009a) may be captured by the local theory. Using this theory as a guide, we then proceed to construct a parameter space (based on the scales associated with the wave and shear flow) for investigation using fully nonlinear direct numerical simulations (DNS). In cases where wave-breaking leads to turbulence, a key question to be addressed is: how do the eventual bulk momentum and energy exchanges between the mean flow and perturbations depend on the route to wave breaking? An obvious point of reference is the vertical critical level problem, where it has been typically observed that a significant fraction of the incoming wave energy is absorbed by the mean flow near the critical level, even in cases where wave-breaking results in considerable turbulent dissipation (Howland et al., 2021; Winters and D’Asaro, 1994; Thorpe, 1981).
@@ 2.1 Initial value problem
The wave Reynolds number $Re_{w}\equiv N^{*}/(K^{*2}\nu^{*})$ (cf. Fringer and Street, 2003) and Prandtl number $Pr\equiv\nu^{*}/\kappa^{*}$ are defined in terms of the molecular kinematic viscosity $\nu^{*}$ and buoyancy diffusivity $\kappa^{*}$. Note that, under this choice of non-dimensionalisation, the coefficient in front of the buoyancy term in the momentum equation (3a) is unity.
@@
Figure 1: Schematic showing a representative initial condition for the problem under consideration. The shape and location of the incident wave packet is shown by plotting isocontours of the $y$-component of the velocity field ($v$). Notably, the wave packet is attenuated in $y$ based on the exponential envelope defined in (7). The tilt of the isocontours in the $xz$ plane is governed by the wavenumber ratio $k_{x}/k_{z}$.
@@
with amplitudes given by (5). Taking $y_{0}<0$, the sign of $k_{y0}$ is chosen to be positive such that the wave packet propagates in the positive $y$-direction. We note that wave packets of this form have also been studied by both Staquet and Huerre (2002) and Bakas and Farrell (2009a). In dimensionless form, the background shear flow into which the wave will propagate is
@@ 2.2 Linear dynamics
where $\epsilon$ is a small parameter characterising the relative scale of phase variations. In the problem at hand, the WKBJ framework describes waves with wavelengths that are asymptotically small relative to the width of the wave envelope $a$ and the width of shear layer $h$. Precisely, a local frequency $\omega(\mathbf{x},t)$ and wavenumber $\mathbf{k}(\mathbf{x},t)$ may be defined in terms of the phase function $\varphi(\mathbf{x},t)$ as
@@
where $\Omega\equiv\omega-Uk_{x}$ is the intrinsic frequency measured by an observer travelling at the speed of the background flow. The frequency $\omega$ measured by a stationary observer is often referred to as the absolute (or extrinsic) frequency. The two are related by the Doppler shift term $U(y)k_{x}$. For a nontrivial balance in (12), we must have $\Omega=O(1)$. It follows that the first two terms are $O(\epsilon^{-2})$, so that the standard dispersion relation (6) is recovered when $(\partial U/\partial y)k_{x}k_{y}=O(\epsilon^{-1})$. Because $k_{x}$ is a priori assumed $O(1/\epsilon)$, the usual approach in the literature is to assume that $U$ (and hence $\partial U/\partial y$) is $O(\epsilon)$ (see, e.g. Olbers, 1981). However, it will be shown below that, due to the fact that the background flow does not vary in the $x$ direction, $k_{x}$ is constant along rays. Therefore, the assumption that $k_{x}=1/\epsilon$ and hence $U=O(\epsilon)$ may be relaxed (provided the wave amplitude varies sufficiently slowly in $x$), as Bretherton (1966) himself pointed out. Thus we may instead consider the case $k_{x}=O(1)$ with $U=O(1)$, in which case local polarisation relations are given by